# Invariants of Finite Orthogonal Groups of Plus Type in Odd Characteristic

### v2

@@ 5.8. Computing the Sylow Invariants
Subducting $\xi_{0}^{q}-x_{1}^{q-1}\xi_{1}$ gives $\psi_{1}(\xi_{0})$, which has lead monomial $y_{2}^{q}x_{2}^{q}$. Subducting $\xi_{1}^{q}-x_{1}^{q-1}\xi_{2}$ gives $\psi_{1}(\xi_{1})$, which has lead monomial $y_{2}^{q^{2}}x_{2}^{q}$. This suggests that a Khovanskii bases for $S_{3}^{P_{3}}$ should include $\psi_{1}(\xi_{0})$ and $\psi_{1}(\xi_{1})$. We know that $S_{2}^{P_{2}}$ is generated by $\xi_{0}$, $\xi_{1}$ and the orbit products of the variables. We also know that this set is a Khovanskii basis. Using the isomorphism from $S_{2}^{P_{2}}$ to $\overline{S}^{P_{2}}$, we conclude that $\{N(y_{2}),N(y_{3}),N(x_{3}),N(x_{2}),\psi_{1}(\xi_{0}),\psi_{1}(\xi_{1})\}$ is a Khovanskii basis for the subalgebra $\overline{S}^{p_{2}}\subset S_{3}^{P_{3}}$. Therefore the tête-a-têtes $\psi_{1}(\xi_{0})^{q}-N(x_{2})^{q-1}\psi_{1}(\xi_{1})$, $\psi_{1}(\xi_{1})^{q}-N(x_{2})^{q}N(y_{2})$ and $\psi_{1}(\xi_{2})-N(x_{2})N(y_{2})^{q}$ subduct to zero. As a result, the tête-a-têtes $(\xi_{0}^{q}-x_{1}^{q-1}\xi_{1})^{q}$, $(\xi_{1}^{q}-x_{1}^{q-1}\xi_{2})^{q}$and $\xi_{2}^{q}-x_{1}^{q-1}\xi_{3}$ subduct to zero. To prove that adjoining $\psi_{1}(\xi_{0})$ and $\psi_{1}(\xi_{1})$ to our generating set gives a Khovanskii basis, we need only show that the tête-a-tête $\xi_{3}^{q}-x_{1}^{q}N(y_{1})$ subducts to zero. We know that this tête-a-tête subducts to zero over the Hook invariants but it is not immediately clear that this subduction can be constructed over the Sylow invariants. It follows from the proof of Theorem 5.9 that $S_{3}^{P_{3}}$ is the free module over the algebra generated by $\mathcal{H}_{0}$ with module generators
$\mathcal{B}_{H}\mathcal{C}=\{\xi_{0}^{a_{0}}\xi_{1}^{a_{1}}\xi_{2}^{a_{2}}\xi_{3}^{a_{3}}\psi_{1}(\xi_{0})^{b_{1}}\psi_{1}(\xi_{1})^{b_{2}}\mid a_{i}<q,\,b_{i}<q\}.$
The lead monomials are given by
$\operatorname{LT}(\mathcal{B}_{H}\mathcal{C})=\{x_{1}^{a_{0}+a_{1}+a_{2}+a_{3}}y_{1}^{a_{0}+a_{1}q+a_{2}q^{2}+a_{3}q^{3}}x_{2}^{q(b_{1}+b_{2})}y_{2}^{qb_{1}+q^{2}b_{2}}\mid a_{i}<q,\,b_{i}<q\}.$
Therefore $\operatorname{LT}(\mathcal{B}_{H}\mathcal{C})$ generates a free module of rank $q^{6}$ over the algebra generated by $\operatorname{LT}(\mathcal{H}_{0})$. This module is a subalgebra of $\operatorname{LT}(S_{3}^{P_{3}})$. By comparing Hilbert series and using the fact that $\operatorname{LT}(S_{3}^{P_{3}})$ and $S_{3}^{P_{3}}$ have the same Hilbert series, we see that $\mathcal{H}_{0}\cup\mathcal{B}_{H}\mathcal{C}$ is a Khovanskii basis for $S_{3}^{P_{3}}$. As a consequence, $S_{3}^{P_{3}}$ is a complete intersection with relations constructed by subducting the tête-a-têtes $\xi_{0}^{q^{2}}-(x_{1}^{q-1}\xi_{1})^{q}$, $\xi_{1}^{q^{2}}-(x_{1}^{q-1}\xi_{2})^{q}$, $\xi_{2}^{q}-x_{1}^{q-1}\xi_{3}$ and $\xi_{3}^{q}-x_{1}^{q}N(y_{1})$. Furthermore, we could replace the module generators $\mathcal{B}_{H}\mathcal{C}$ with $\{\xi_{0}^{e_{0}}\xi_{1}^{e_{1}}\xi_{2}^{e_{2}}\xi_{3}^{e_{3}}\mid e_{0},e_{1}<q^{2},\,e_{2},e_{3}<q\}$.
@@
Applying Lemma 5.11 inductively shows that the elements of $\mathcal{B}_{H}\mathcal{C}$ are of the form $\prod\psi_{j}(\xi_{i})^{e_{ij}}$ with $e_{ij}<q$. Using Lemma 5.10, careful book-keeping shows that $\operatorname{LT}(\mathcal{B}_{H}\mathcal{C})$ has the same cardinality as $\mathcal{B}_{H}\mathcal{C}$ and that $\operatorname{LT}(\mathcal{B}_{H}\mathcal{C})$ generates a free module over the algebra generated by $\operatorname{LT}(\mathcal{H}_{0})$. This module is a subalgebra of $\operatorname{LT}(S_{m}^{P_{m}})$. By comparing Hilbert series and using the fact that $\operatorname{LT}(S_{m}^{P_{m}})$ and $S_{m}^{P_{m}}$ have the same Hilbert series, we see that $\mathcal{H}_{0}\cup\mathcal{B}_{H}\mathcal{C}$ is a Khovanskii basis for $S_{m}^{P_{m}}$. Let $\widetilde{\mathcal{B}_{H}}$ denote the subset of $\mathcal{B}_{H}\mathcal{C}$ consisting of elements of the form $\psi_{j}(\xi_{i})$. Since the algebra generated by $\operatorname{LT}(\mathcal{H}_{0}\cup\widetilde{\mathcal{B}_{H}})$ coincides with the algebra generated by $\operatorname{LT}(\mathcal{H}_{0}\cup\mathcal{B}_{H}\mathcal{C})$, we see that $\mathcal{H}_{0}\cup\widetilde{\mathcal{B}_{H}}$ is also a Khovanskii basis for $S_{m}^{P_{m}}$.
Note that ~$\widetilde{\mathcal{B}}_{H}=\{\psi_{j}(\xi_{i})\mid 0\leq i\leq n-3-2j\}$. The non-trivial tête-a-têtes are $\psi_{j}(\xi_{i})^{q}-\psi_{j}(\xi_{i+1})N(x_{j+1})^{q-1}$ for $i<n-3-2j$ and $\psi_{j}(\xi_{i})^{q}-N(y_{j+1})N(x_{j+1})^{q}$ for $i=n-3-2j$. Since $\mathcal{H}_{0}\cup\widetilde{\mathcal{B}_{H}}$ is a Khovanskii basis, each of these tête-a-têtes subducts to zero. The resulting relations are independent and generate the ideal of relations. The number of generators minus the number of independent relations is $n$, proving that $S_{m}^{P_{m}}$ is a complete intersection. ∎
@@
It follows from the proof of Theorem 5.13 that $\mathcal{B}_{H}\mathcal{C}$ consists of the monomial factors of $\prod\{\psi_{j}(\xi_{i})^{q-1}\mid 0\leq i\leq n-3-2j\}$.
@@
$\Gamma_{0}=\prod_{i=0}^{n-3}\xi_{i}^{e_{i}}\text{ where }e_{i}=q^{m-1-\lfloor i/2\rfloor}-1\ .$
@@
Using Theorem 5.13 and its proof, each of the tête-a-têtes $\xi_{n-3}^{q}-x_{1}^{q}N(y_{1})$, $\xi_{n-4}^{q}-x_{1}^{q-1}\xi_{n-3}$, $\xi_{n-5}^{q^{2}}-x_{1}^{q-1}\xi_{n-4}^{q}$, $\xi_{n-6}^{q^{2}}-x_{1}^{q-1}\xi_{n-5}^{q},\dots,\xi_{1}^{q^{m-1}}-x_{1}^{q-1}\xi_{2}^{q^{m-2}}$, and $\xi_{0}^{q^{m-1}}-x_{1}^{q-1}\xi_{1}^{q^{m-2}}$ subducts to zero. Applying Theorem 5.9 and comparing the number of monomial factors of $\Gamma_{0}$ with the number of module generators for $S_{m}^{P_{m}}$ over $\mathbb{F}_{q}[\mathcal{H}_{0}]$ yields the result. ∎
@@
$S_{m}^{P_{m}}$ is minimally generated by $\mathcal{H}_{0}\cup\{\xi_{0},\xi_{1},\dots,\xi_{n-3}\}$.
@@
$\mathcal{F}(S_{m}^{P_{m}})=\mathbb{F}_{q}(N(y_{m}),N(x_{1}),N(x_{2}),\dots,N(x_{m}),\xi_{0},\xi_{1},\dots,\xi_{m-2})$
@@
$\mathcal{V}(N(y_{m}),N(x_{1}),N(x_{2}),\dots,N(x_{m}),\xi_{0},\xi_{1},\dots,\xi_{m-2})=\operatorname{Span}_{\overline{\mathbb{F}_{q}}}\{e_{1},e_{2},\dots,e_{m-1}\}.$
The remaining $\xi_{i}$ all vanish on this set and so to cut this variety down to a point (the origin) we need to add in the remaining $m-1$ norms. It remains to show that $\xi_{m-1},\xi_{m},\dots,\xi_{n-3}$ are also necessary as generators.
Otherwise, one of these $\xi_{i}$ can be expressed as a polynomial $f$ in the other $32n-$ 3 generators. Since $\operatorname{LT}(\xi_{i})=y_{1}^{q^{i}}x_{1}$ this must also be the lead term of $f$ as an element of $S_{m}$. This lead term cannot be expressed as a product of the lead terms of the other $32m-$ 3 generators and thus it must arise from a tête-a-tête. This tête-a-tête must be in degree $q^{i}+1$ and involve at least one generator whose lead term is greater than $y_{1}^{q^{i}}x_{1}$. But any of these $32m-$ 3 generators whose lead term is greater than $y_{1}^{q^{i}}x_{1}$ also has degree greater than $q^{i}+1$. Hence no such tête-a-tête can exist. ∎
@@ 5.18. Computing the Borel Invariants
Let $B$ denote the Borel subgroup of upper triangular matrices. The Sylow subgroup $P$ is a normal subgroup of $B$ and we identify $B/P$ with the subgroup of diagonal matrices. Let $\mathcal{H}_{0}$ denote the homogeneous system of parameters for $S^{P}$ given by the orbit products over $P$ of the variables and let $\mathbb{F}_{q}[\mathcal{H}_{0}]$ denote the algebra generated by $\mathcal{H}_{0}$. Define $\mathcal{B}_{H}$ and $\mathcal{C}$ as in the proof of Theorem 5.9. Then
@@
$N(w)b$ =($=\left(\prod\{wg\mid g\in P\}\right)b=\prod\{wgb\mid g\in P\}=\prod\{wb\overline{g}\mid g\in P\}$
=$=\prod\{cw\overline{g}\mid g\in P\}=c^{q^{\ell}}\prod\{wg\mid g\in P\}=c^{q^{\ell}}N(w)$
@@
Since the elements of $\mathcal{H}_{0}$ are orbit products over $P$, it follows from Lemma 5.19 that $B/P$ acts on $\mathbb{F}_{q}[\mathcal{H}_{0}]$. Arguing as in the proof of Theorem 5.13, the elements of $\mathcal{B}_{H}\mathcal{C}$ are of the form $\psi_{j}(\xi_{i})$. Since $\psi_{j}(\xi_{i})$ is Borel invariant, we have
@@
The proof follows from the computation of $\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P}$. It follows from Lemma 5.19 that $N(y_{i})^{q-1}$ and $N(x_{i})^{q-1}$ are Borel invariant. Furthermore, $B$ contains the linear transformation which scales $y_{i}$ by $c\in\mathbb{F}_{q}$, takes $x_{i}$ to $c^{-1}x_{i}$, and fixes the other variables. Therefore no smaller power of $N(y_{i})$ or $N(x_{i})$ is Borel invariant. Any element of $B/P$ which scales $y_{i}$ by $c$ must scale $x_{i}$ by $c^{-1}$. Therefore $N(y_{i})N(x_{i})$ is Borel invariant. It follows from this that $\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P}$ is the complete intersection generated by $\widetilde{\mathcal{H}_{0}}$ subject to the relations $(N(y_{i})N(x_{i}))^{q-1}=N(y_{i})^{q-1}N(x_{i})^{q-1}$ and $\widetilde{\mathcal{H}_{0}}$ and is a Khovanskii basis for $\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P}$. Since there are no new tête-a-têtes involving $\mathcal{B}_{H}\mathcal{C}$, $\widetilde{\mathcal{H}_{0}}\cup\mathcal{B}_{H}\mathcal{C}$ is a Khovanskii basis for $S^{B}$, as required. ∎
The Weyl group, $W$, of ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ is isomorphic to the wreath product $C_{2}\wr\Sigma_{m}$. We can identify $W$ with a subgroup of ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$: $\Sigma_{m}$ acts by permuting the hyperbolic pairs and the additional generator, say $\omega$, interchanges $y_{m}$ and $x_{m}$ while fixing the other variables.
@@
$[G:B]=2\frac{q^{m}-1}{q-1}\prod_{j=1}^{m-1}\frac{q^{2j}-1}{q-1}\equiv_{(p)}2$
@@
To compute the Reynolds operator we need to choose left coset representatives for $B$ in ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$. We use the double coset partition to do this. The double cosets, $BwB$ for $w\in W$, partition ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ so we can choose coset representatives of the form $wb$ for $w\in W$ and $b\in B$. Observe that the number of left cosets in the double coset $BwB$ is given by the index $[B:B\cap w^{-1}Bw]=[P:P\cap w^{-1}Pw]$, which divides $q^{m(m-1)}$.
@@
$\operatorname{LT}\left(\mathcal{R}(N(y_{1})^{q-1})\right)=y_{1}^{(q-1)q^{n-2}}.$

### v3

@@ 5.10. Computing the Sylow Invariants
Subducting $\xi_{1}^{q}-x_{3}^{q-1}\xi_{2}$ using Equation 4 gives $\psi_{[3,1]}(\xi_{1})$, which has lead monomial $y_{2}^{q}x_{2}^{q}$. Adding $\psi_{[3,1]}(\xi_{1})$ to the generating set and subducting $\xi_{2}^{q}-x_{3}^{q-1}\xi_{3}$ using Lemma 5.2 gives $\psi_{[3,1]}(\xi_{2})$, which has lead monomial $y_{2}^{q^{2}}x_{2}^{q}$. This suggests that a Khovanskii bases for $S_{3}^{P_{3}}$ should include $\psi_{[3,1]}(\xi_{1})$ and $\psi_{[3,1]}(\xi_{2})$. We know that $S_{2}^{P_{2}}$ is generated by $\xi_{1}$, $\xi_{2}$ and the orbit products of the variables (Theorem 5.5). We also know that this set is a Khovanskii basis (Lemma 5.4). Recall that $\overline{S}=\mathbb{F}_{q}[Y_{2},Y_{1},X_{1},X_{2}]$. Using the isomorphism from $S_{2}^{P_{2}}$ to $\overline{S}^{P_{2}}$, we conclude that
$\{N(y_{2}),N(y_{1}),N(x_{1}),N(x_{2}),\psi_{[3,1]}(\xi_{1}),\psi_{[3,1]}(\xi_{2})\}$
is a Khovanskii basis for the subalgebra $\overline{S}^{P_{2}}\subset S_{3}^{P_{3}}$. Therefore the tête-a-têtes $\psi_{[3,1]}(\xi_{1})^{q}-N(x_{2})^{q-1}\psi_{[3,1]}(\xi_{2})$ and $\psi_{[3,1]}(\xi_{2})^{q}-N(x_{2})^{q}N(y_{2})$ subduct to zero in $\overline{S}^{P_{2}}$ (these subduction are also subductions in $S_{3}^{P_{3}}$) . Since $\psi_{[3,1]}(\xi_{3})-N(x_{2})N(y_{2})^{q}$ subducts to zero in $\overline{S}^{P_{2}}$, $\psi_{[3,1]}(\xi_{3})$ is not needed in the Khovanskii basis and $\xi_{3}^{q}-x_{3}^{q-1}\xi_{4}$ subducts to zero in $S_{3}^{P_{3}}$. To prove that adjoining $\psi_{[3,1]}(\xi_{1})$ and $\psi_{[3,1]}(\xi_{2})$ to our generating set gives a Khovanskii basis, we need only show that the tête-a-tête $\xi_{4}^{q}-x_{3}^{q}N(y_{3})$ subducts to zero. We know that this tête-a-tête subducts to zero over the Hook invariants but it is not immediately clear that this subduction can be constructed over the Sylow invariants. It follows from the proof of Theorem 5.11 that $S_{3}^{P_{3}}$ is the free module over the algebra generated by $\mathcal{H}_{0}$ with module generators
$\mathcal{B}_{H}\mathcal{C}=\{\xi_{1}^{a_{1}}\xi_{2}^{a_{1}}\xi_{3}^{a_{3}}\xi_{4}^{a_{4}}\psi_{[3,1]}(\xi_{1})^{b_{1}}\psi_{[3,1]}(\xi_{2})^{b_{2}}\mid a_{i}<q,\,b_{i}<q\}.$
Therefore the Hilbert series of $S_{3}^{P_{3}}$, and its lead term algebra, is given by
$HS(S_{3}^{P_{3}},t)=HS(\mathcal{H}_{0},t)\sum_{\gamma\in\mathcal{B}_{H}\mathcal{C}}t^{\deg(\gamma)}$
where
$HS(\mathcal{H}_{0},t)=HS(\operatorname{LT}(\mathcal{H}_{0}),t)=\prod_{i=1}^{3}\frac{1}{(1-t^{q^{i-1}})(1-t^{q^{5-i}})}\ .$
The lead monomials of $\mathcal{B}_{H}\mathcal{C}$ are given by
$\operatorname{LM}(\mathcal{B}_{H}\mathcal{C})=\{x_{3}^{a_{1}+a_{2}+a_{3}+a_{4}}y_{3}^{a_{1}+a_{2}q+a_{3}q^{2}+a_{4}q^{3}}x_{2}^{q(b_{1}+b_{2})}y_{2}^{qb_{1}+q^{2}b_{2}}\mid a_{i}<q,\,b_{i}<q\}.$
Therefore $\operatorname{LM}(\mathcal{B}_{H}\mathcal{C})$ generates a free module of rank $q^{6}$ over $\operatorname{LT}(\mathcal{H}_{0})$. The Hilbert series of the vector space span of $\operatorname{LM}(\mathcal{B}_{H}\mathcal{C})$ is
$HS(\operatorname{LT}(\mathcal{B}_{H}\mathcal{C},t))$ $=\Big(\sum_{b_{1}=0}^{q-1}t^{2b_{1}q}\Big)\Big(\sum_{b_{2}=0}^{q-1}t^{b_{2}(q+q^{2})}\Big)\prod_{i=1}^{4}\Bigg(\sum_{a_{i}=0}^{q-1}t^{a_{i}(q^{i-1}+1)}\Bigg)$
$=\Big(\frac{1-t^{2q^{2}}}{1-t^{2q}}\Big)\Big(\frac{1-t^{q^{2}+q^{3}}}{1-t^{q+q^{2}}}\Big)\prod_{i=1}^{4}\Bigg(\frac{1-t^{q^{i}+q}}{1-t^{q^{i-1}+1}}\Bigg)$
$=\sum_{\gamma\in\mathcal{B}_{H}\mathcal{C}}t^{\deg(\gamma)}\ .$
The module generated by $\operatorname{LM}(\mathcal{B}_{H}\mathcal{C})$ over $\operatorname{LT}(\mathcal{H}_{0})$ is a submodule of $\operatorname{LT}(S_{3}^{P_{3}})$. By the above this module has the same Hilbert series as $\operatorname{LT}(S_{3}^{P_{3}})$, and thus they are equal. Therefore $\mathcal{H}_{0}\cup\mathcal{B}_{H}\mathcal{C}$ is a Khovanskii basis for $S_{3}^{P_{3}}$. Define
$\widetilde{\mathcal{B}_{H}}:=\{\xi_{1},\xi_{2},\xi_{3},\xi_{4},\psi_{[3,1]}(\xi_{1}),\psi_{[3,1]}(\xi_{2})\}\ .$
[…]
$\psi_{[3,1]}(\xi_{1})=\xi_{1}^{q}-\xi_{2}x_{3}^{q-1}+\xi_{1}x_{3}^{2q-2}$
and
$\psi_{[3,1]}(\xi_{2})=\xi_{2}^{q}-\xi_{3}x_{3}^{q-1}-\xi_{1}^{q}x_{3}^{q^{2}-q}+\xi_{1}x_{3}^{q^{2}+q-2}-\psi_{[3,1]}(\xi_{1})x_{3}^{q^{2}-q}\ .$
Substituting these expressions for $\psi_{[3,1]}(\xi_{1})$ and $\psi_{[3,1]}(\xi_{2})$ into the subductions of the tête-a-têtes $\xi_{3}^{q}-x_{3}^{q-1}\xi_{4}$ and $\xi_{4}^{q}-x_{3}^{q}N(y_{1})$ gives relations which rewrite $\xi_{3}^{q}$ and $\xi_{4}^{q}$. Substituting the expressions for $\psi_{[3,1]}(\xi_{1})$ and $\psi_{[3,1]}(\xi_{2})$ into the relations associated to the last two tête-a-têtes gives relations which rewrite $\xi_{1}^{q^{2}}$ and $\xi_{2}^{q^{2}}$ Therefore, using the smaller generating set, we can replace the module generators $\mathcal{B}_{H}\mathcal{C}$ with $\{\xi_{1}^{e_{1}}\xi_{2}^{e_{2}}\xi_{3}^{e_{3}}\xi_{4}^{e_{4}}\mid e_{1},e_{2}<q^{2},\,e_{3},e_{4}<q\}$.
@@
[…]
Note that $\widetilde{\mathcal{B}}_{H}=\{\psi_{[m,j]}(\xi_{i})\mid 0\leq j\leq m-2\text{ and }1\leq i\leq n-2-2j\}$. The non-trivial tête-a-têtes are $\psi_{[m,j]}(\xi_{i})^{q}-\psi_{[m,j]}(\xi_{i+1})N(x_{j+1})^{q-1}$ for i<n22ji<n-2-2j and $\psi_{[m,j]}(\xi_{i})^{q}-N(y_{j+1})N(x_{j+1})^{q}$ for i=n22ji=n-2-2j. Since $\mathcal{H}_{0}\cup\widetilde{\mathcal{B}_{H}}$ is a Khovanskii basis, each of these tête-a-têtes subducts to zero. The resulting relations are independent and generate the ideal of relations. The number of generators minus the number of independent relations is $n$, proving that $S_{m}^{P_{m}}$ is a complete intersection. ∎
@@
It follows from the proof of Theorem 5.15 that $\mathcal{B}_{H}\mathcal{C}$ consists of the monomial factors of $\prod\{\psi_{[m,j]}(\xi_{i})^{q-1}\mid 0\leq j\leq m-2\text{ and }1\leq i\leq n-2-2j\}$.
@@
$\Gamma_{0}=\prod_{i=1}^{n-2}\xi_{i}^{e_{i}}\text{ where }e_{i}=q^{m-\lceil i/2\rceil}-1\ .$
@@
Using Theorem 5.15 and its proof, each of the tête-a-têtes $\xi_{n-2}^{q}-x_{m}^{q}N(y_{m})$, $\xi_{n-3}^{q}-x_{m}^{q-1}\xi_{n-2}$, $\xi_{n-4}^{q^{2}}-x_{m}^{q(q-1)}\xi_{n-3}^{q}$, $\xi_{n-5}^{q^{2}}-x_{m}^{q(q-1)}\xi_{n-4}^{q},\dots,\xi_{2}^{q^{m-1}}-x_{m}^{q^{m-2}(q-1)}\xi_{3}^{q^{m-2}}$, and $\xi_{1}^{q^{m-1}}-x_{m}^{q^{m-2}(q-1)}\xi_{2}^{q^{m-2}}$ subducts to zero. Applying Theorem 5.11 and comparing the number of monomial factors of $\Gamma_{0}$ with the number of module generators for $S_{m}^{P_{m}}$ over $\mathbb{F}_{q}[\mathcal{H}_{0}]$ yields the result. ∎
@@
$S_{m}^{P_{m}}$ is minimally generated by $\mathcal{H}_{0}\cup\{\xi_{1},\xi_{2},\dots,\xi_{n-2}\}$.
@@
$\mathcal{F}(S_{m}^{P_{m}})=\mathbb{F}_{q}(N(y_{1}),N(x_{1}),N(x_{2}),\dots,N(x_{m}),\xi_{1},\xi_{2},\dots,\xi_{m-1})$
@@
$\mathcal{V}(N(y_{1}),N(x_{1}),N(x_{2}),\dots,N(x_{m}),\xi_{1},\xi_{2},\dots,\xi_{m-1})=\operatorname{Span}_{\overline{\mathbb{F}}_{q}}\{e_{2},e_{2},\dots,e_{m}\}.$
The remaining $\xi_{i}$ all vanish on this set and so to cut this variety down to a point (the origin) we need to add in the remaining m1m-1 norms. It remains to show that $\xi_{m},\xi_{m+1},\dots,\xi_{n-2}$ are also necessary as generators.
Otherwise, one of these $\xi_{i}$ can be expressed as a polynomial $f$ in the other 2n32n-3 generators. Using the lex order, since $\operatorname{LT}(\xi_{i})=y_{m}^{q^{i-1}}x_{m}$, this must also be the lead term of $f$ as an element of $S_{m}$. This lead term cannot be expressed as a product of the lead terms of the other 2m32m-3 generators and thus it must arise from a tête-a-tête. This tête-a-tête must be in degree $q^{i-1}+1$ and involve at least one generator whose lead term is greater than $y_{m}^{q^{i-1}}x_{m}$. But any of these 2m32m-3 generators whose lead term is greater than $y_{m}^{q^{i-1}}x_{m}$ also has degree greater than $q^{i-1}+1$. Hence no such tête-a-tête can exist. ∎
@@ 5.20. Computing the Borel Invariants
Let $B$ denote the Borel subgroup of upper triangular matrices in $\text{O}^{+}_{{2m}}(\mathbb{F}_{q})$}}. The Sylow subgroup $P$ is a normal subgroup of $B$ and we identify $B/P$ with the subgroup of diagonal matrices, which we denote by $T$. Let $N$ denote the normaliser of $T$ in $\text{O}^{+}_{{2m}}(\mathbb{F}_{q})$}}. The Weyl group, $W:=N/T$, is isomorphic to the wreath product $C_{2}\wr\Sigma_{m}$. We can identify $W$ with a subgroup of $\text{O}^{+}_{{2m}}(\mathbb{F}_{q})$}}: the group $\Sigma_{m}$ acts by permuting the hyperbolic pairs and the additional generator, say $\omega$, interchanges $y_{1}$ and $x_{1}$ while fixing the other variables. Note that $B$ and $N$ are not a $BN$-pair for $\text{O}^{+}_{{2m}}(\mathbb{F}_{q})$}}. Most of the axioms are satisfied and ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}=BNB$ but $\omega B\omega=B$ (see [27, p. 85] and [16, §1.7.5]). The subgroup $\text{SO}^{+}_{{2m}}(\mathbb{F}_{q})$}} contains $B$ and has a $BN$-pair with Weyl $W^{\prime}$, an index $2$ subgroup of $W$. The Dynkin type of $W^{\prime}$ is $D_{m}$ (see [16, §1.7.9, §1.3.16]).
Let $\mathcal{H}_{0}$ denote the homogeneous system of parameters for $S^{P}$ given by the orbit products over $P$ of the variables and let $\mathbb{F}_{q}[\mathcal{H}_{0}]$ denote the algebra generated by $\mathcal{H}_{0}$. Define $\mathcal{B}_{H}$ and $\mathcal{C}$ as in the proof of Theorem 5.11. Then
@@
$N(w)b$ $=\left(\prod\{wg\mid g\in P\}\right)b=\prod\{wgb\mid g\in P\}=\prod\{wb\overline{g}\mid g\in P\}$
$=\prod\{cw\overline{g}\mid g\in P\}=c^{q^{\ell}}\prod\{wg\mid g\in P\}=c^{q^{\ell}}N(w)$
@@
Since the elements of $\mathcal{H}_{0}$ are orbit products over $P$, it follows from Lemma 5.21 that $B/P$ acts on $\mathbb{F}_{q}[\mathcal{H}_{0}]$. Arguing as in the proof of Theorem 5.15, the elements of $\mathcal{B}_{H}\mathcal{C}$ are monomials in $\psi_{[m,j]}(\xi_{i})$. Since $\psi_{[m,j]}(\xi_{i})$ is Borel invariant, we have
@@
The proof follows from the computation of $\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P}$. From Lemma 5.21, $N(y_{i})^{q-1}$ and $N(x_{i})^{q-1}$ are Borel invariant. Furthermore, $B$ contains the linear transformation which scales $y_{i}$ by $c\in\mathbb{F}_{q}$, takes $x_{i}$ to $c^{-1}x_{i}$, and fixes the other variables. Therefore no smaller power of $N(y_{i})$ or $N(x_{i})$ is Borel invariant. Any element of $B/P$ which scales $y_{i}$ by $c$ must scale $x_{i}$ by $c^{-1}$. Therefore $N(y_{i})N(x_{i})$ is Borel invariant. It follows from this that $\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P}$ is the complete intersection generated by $\widetilde{\mathcal{H}_{0}}$ subject to the relations $(N(y_{i})N(x_{i}))^{q-1}=N(y_{i})^{q-1}N(x_{i})^{q-1}$ and $\widetilde{\mathcal{H}_{0}}$ and is a Khovanskii basis for $\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P}$. Let $\widetilde{K}$ denote the algebra generated by the lead monomials of $\widetilde{\mathcal{H}_{0}}$. Since $\widetilde{\mathcal{H}_{0}}$ is a Khovanskii basis for $\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P}$, the Hilbert series of $\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P}$ equals the Hilbert series of $\widetilde{K}$. Furthermore
$HS(S^{B},t)=HS(\mathbb{F}_{q}[\mathcal{H}_{0}]^{B/P},t)\left(\sum_{\gamma\in\mathcal{B}_{H}\mathcal{C}}t^{\deg(\gamma)}\right)\ .$
Therefore
$HS(\operatorname{LT}(S^{B}),t)=HS(S^{B},t)=HS(\widetilde{K},t)\left(\sum_{\gamma\in\mathcal{B}_{H}\mathcal{C}}t^{\deg(\gamma)}\right)$
which proves that $\widetilde{\mathcal{H}_{0}}\cup\mathcal{B}_{H}\mathcal{C}$ is a Khovanskii basis for $S^{B}$. This means that the non-trivial tête-a-têtes from $\mathcal{B}_{H}\mathcal{C}$ subduct to zero in $S^{B}$. These subductions, along with the relations coming from $\widetilde{\mathcal{H}_{0}}$, give a generating set for the ideal of relations, proving that $S^{B}$ is a complete intersection. ∎
@@
$G:B]=2\frac{q^{m}-1}{q-1}\prod_{j=1}^{m-1}\frac{q^{2j}-1}{q-1}\equiv_{(p)}2$
@@
To compute the Reynolds operator we need to choose left coset representatives for $B$ in $\text{O}^{+}_{{2m}}(\mathbb{F}_{q})$}}. We use the double coset partition to do this. The double cosets, $BwB$ for $w\in W$, partition $\text{O}^{+}_{{2m}}(\mathbb{F}_{q})$}} so we can choose coset representatives of the form $wb$ for $w\in W$ and $b\in B$. Observe that the number of left cosets in the double coset $BwB$ is given by the index $B:B\cap w^{-1}Bw]=[P:P\cap w^{-1}Pw]$, which divides $q^{m(m-1)}$.
@@
$\operatorname{LT}\left(\mathcal{R}(N(y_{m})^{q-1})\right)=y_{m}^{(q-1)q^{n-2}}.$